OpenStudy (anonymous):

a man deposits a fixed amount of money in the beginning of every month in a bank which gives 10% simple interest per year , if the total interest in the end of a year is 117 , calculate the monthly deposit

OpenStudy (ash2326):

@hossam do you have the answer?

OpenStudy (anonymous):

yes its 180

OpenStudy (ash2326):

Okay, I thought something else let me try again

OpenStudy (anonymous):

okay !!!!!

OpenStudy (ash2326):

Bank will give interest for deposit depending on time, His each deposit will earn a different interest. His first deposit will earn interest for 12 months or 12/12 years His second deposit will earn interest for 11 months or 11/12 years His third deposit will earn interest for 10 months similarly His twelfth deposit will earn interest for 1 month or 1/12 years Let his each deposit be P so first deposit will earn interest i1 \[i1=P \times R \% \times t\] or \[i1= \frac{P \times R \times 12}{100 \times 12}\] R=10 so we get \[i1=\frac{P \times 12}{10 \times 12}\] for second deposit \[i2=\frac{P \times 11}{10 \times 12}\] for third deposit \[i2=\frac{P \times 10}{10 \times 12}\] similarly we can write for each month for last deposit \[i12=\frac{P \times 1}{10 \times 12}\] so total interest \[i=i1+i2+i3+....+i12\] we are given the total interest is 117 so \[117=\frac{P \times 12}{10 \times 12}+\frac{P \times 11}{10 \times 12}+........+\frac{P \times 1}{10 \times 12}\] I have written only few terms above, let's take common term out on right hand side \[117=\frac{P}{10 \times 12}(12+11+10+9+8+7+6+5+4+3+2+1)\] we get now \[117=\frac{P}{10 \times 12}(78)\] \[P= \frac{117 \times 10 \times 12}{78}\] so we get \[P=1.5 \times 10 \times 12= 180\] so his each deposit is 180

OpenStudy (ash2326):

Did you understand?

OpenStudy (anonymous):

thanks a million , yes i did